How many grams of Ag₂ CrO4 will precipitate when 86ml of 0.2 molar Ag₂Cro, is added to 500ml of 6.3 molar K₂ CrO4? Molar mass of Ag₂ CrO4 = 331.74 g/mol







Balanced Equation: 2AgNO3 + K₂CrO4 Ag₂ CrO4 + 2KNO3

1
Expert's answer
2022-03-10T10:47:46-0500

Moles of AgNO3 = 0.2×86÷1000=0.0172

Moles of K2CrO4=6.3×500÷1000= 3.15

Limiting reagent is AgNO3

Moles of Ag2CrO4 formed = 0.0172÷2= 0.0086

Mass of Ag2CrO4 produces = 0.0086×331.74 = 2.85296g